OpenStudy (anonymous):

Find g(1) if g(x) = x2 + 1. 2 3 4

OpenStudy (anonymous):

2

OpenStudy (anonymous):

David to solve these problems - just plug into the equation the value in the parenthesis for x

OpenStudy (unklerhaukus):

\[g(x)=x^2+1\] \[g(1)=(1)^2+1\]\[~~~~~~~~=1+1\]\[~~~~~~~~~=2\]